GlueX TOF Meeting, July 12, 2016
GlueX Time-of-Flight Meeting
Tuesday, July 12, 2016
11:00 am EDT
JLab: CEBAF Center, Room F326/327
- Review of minutes from the previous meeting
- PMT Gains
- Calibration Status
- NIM paper: svn co https://halldsvn.jlab.org/repos/trunk/publications/GlueXTOF
- The BlueJeans Meeting number is 350 531 998.
- Enter BlueJeans Meeting
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .
- JLab: Eugene Chudakov, Mark Ito (chair), Simon Taylor, Beni Zihlmann
There is a recording of this meeting on the BlueJeans site.
Beni reported on tests he has done with the laser.
- Used a spare TOF tube.
- At 1200 V, adjusted laser to give a 150 mV peak pulse. Can increase rate to 10 MHz with no reduction in gain.
- At 1300 V, 150 mV pulse again. At 1 MHz start to see noise (extraneous pulses with low amplitude out of time). These could be due to the PMT heating up.
- At 1400 V
- 150 mV pulse: can go to 10 MHz without loss of amplitude.
- 300 mV pulse: see loss in amplitude at 5 MHz.
- 1.5 V pulse: start seeing loss somewhere between 1 and 2 MHz.
- Behavior consistent with what was observed during the run in the difference in PMT gain between PARA and PERP for some runs. Rates were around 1 MHz and gains varied significantly when intensity changed with the polarization orientation.
- Over the weekend Paul Eugenio, on his way through Newport News, dropped of an 8-stage tube in Beni's office.
- At 1400 V, 150 mV pulse, start to see noise at 2 MHz and a drop in amplitude at 10 MHz.
- At 1250 V, 150 mV pulse, less noise is seen at the higher rates and no drop in amplitude was seen.
- Beni hopes to reduce gain by factor of 4 or 5.
- HV settings already changed to 1250 V for all channels in Hall (from the various values around 1500 V we were using before).
- On the scope he sees that the largest pulses of several Volts are reduced to about 700 mV, a factor of 3 or 4.
- Eugene suggested we measure the gain curve (amplitude as a function of voltage for a fixed input light pulse). The can be used to estimate settings going forward.
- Eugene pointed out that even with this reduction in voltage, when we go to high luminosity, we will right back in the same situation we were in during the last run. Some kind of action is necessary long-term.
- The dynamic range on the ADCs needs to be changed from 2 V full scale to 0.5 V.
- Beni wants to take some cosmics in the new configuration. These data can be compared with those that Sasha Ostrovidov took before the last run.
- Eugene reminded us that we will have to study the effect of these changes on timing.
- Beni found his old wiki page where he studied rates. In that he estimated that we should be aiming for 150 mV pulses from the center of the paddles if we want to have acceptable tube lifetimes.
- A picture is emerging where the amplitude drop is not simply a function of current but rather depends on the PMT gain.
- We need to explore options for replacing the bases, tubes, in all or part of the detector. Base modification would include an amplifier like that used on the tagger hodoscope. Beni will talk to Fernando Barbosa about options.
- We will also have to consult with Tim Whitlatch on the mechanical work in the Hall. We will likely have to dismount the TOF from the FCAL frame to affect the changes. This could be done after the Spring run of 2017. There may be some schedule interference with DIRC installation during that time.
- Mark contacted Ardavan Ghassemi of Hamamatsu for quotes on new tubes. Ardavan is traveling but should be back soon.